OpenStudy (anonymous):

How do I find the numbers that must be excluded from the domain or the rational expression?:

OpenStudy (anonymous):

\[\frac{ 9 }{ x-6 }\]

OpenStudy (anonymous):

domain are those values for which the expression is defined

OpenStudy (anonymous):

here at x=6 the expression becomes undefined thus 6 must be excluded from domain

OpenStudy (anonymous):

How does that work though?

OpenStudy (anonymous):

since the denominator will become zero at x-6=0 i.e. at 6

OpenStudy (anonymous):

So what would you do for \[\frac{ 4 }{ x+3 }\]

OpenStudy (anonymous):

I have a whole section on this and it just gets harder, and I truly don't understand it.

OpenStudy (anonymous):

x + 3 = 0 i.e. at -3

OpenStudy (anonymous):

God, this makes no sense....how is it -3?

OpenStudy (anonymous):

i think u r not getting my point.. oK

OpenStudy (anonymous):

Sorry :/

OpenStudy (anonymous):

Really? I need helppppppppp!

OpenStudy (anonymous):

what is the definition of domain?

OpenStudy (anonymous):

values for which the expression is defined

OpenStudy (anonymous):

and whats the question?

OpenStudy (anonymous):

u need to find those values which are excluded from domain..is it..

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so make that attempt in such a way that the expression becomes undefined

OpenStudy (anonymous):

undefined?

OpenStudy (anonymous):

u don;t understand undefined

OpenStudy (anonymous):

the opposite of defined

OpenStudy (anonymous):

What does that mean, though? what defines a rational expression?

OpenStudy (anonymous):

if u get a real number it becomes defined if u get 1/0 case or 0/0 case it is known as undefined

OpenStudy (anonymous):

So what exactly makes the 6 in the 9/x-6 expression excluded from the domain?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank GOD .. lol

OpenStudy (anonymous):

....what?

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

have u understood now?

OpenStudy (anonymous):

No, otherwise I wouldn't have asked that

OpenStudy (anonymous):

what u haven't understood ?

OpenStudy (anonymous):

Why the 6 is excluded

OpenStudy (anonymous):

because at 6 the expression becomes 9/0 case which is undefined

OpenStudy (anonymous):

what makes it defined?

OpenStudy (anonymous):

if the number in denominator is 0 in any case it is undefined

OpenStudy (anonymous):

otherwise it is defined

OpenStudy (anonymous):

so what about with 4/x+3...why is -3 excluded? how did you get -3?

OpenStudy (anonymous):

since at -3 denominator will become zero

OpenStudy (anonymous):

ohhhhhhh!!! Alright, I see now! What about with more complex expressions like: \[\frac{ x-3 }{ x ^{2}-64 }\]

OpenStudy (anonymous):

u try first

OpenStudy (anonymous):

well, do I have to do anything with the x^2?

OpenStudy (anonymous):

at what value will make denominator zero?

OpenStudy (anonymous):

64?

OpenStudy (anonymous):

make denominator equal to zero x^2-64=0

OpenStudy (anonymous):

find x

OpenStudy (anonymous):

But how? Like: (x+8)(x-8)?

OpenStudy (anonymous):

yeah ... u will get x=8,-8

OpenStudy (anonymous):

So, those are the only answers?